# PhoenixOS: Concurrent OS-level GPU Checkpoint and Restore with Validated Speculation

### v2

@@ Introduction
  3. 3.
    Issue computation kernels with well-defined semantics, e.g., cublasSgemm(...,A,B,...,C,...). They perform a computation and update the result buffers (C), whose read and write semantics are know in the specifications [52].
  4. 4.
    Issue opaque computation kernels, e.g., cudaLaunchKernel (func, ... , args, ...) . They launch kernels written by the developers or just-in-time compiled during runtime, so the reads and writes are unknown to the OS.
@@
Figure 5: An illustration of the GPU kernel [25] written by the user (a) and how PhOS speculates the accessed buffers (b).
Opaque kernels are trickier because the OS only knows the kernel binary and the arguments invoking it. Our observation is that each GPU kernel represents a clear and straightforward computational purpose whose accessed data addresses are directly encoded in the OS-known arguments calling the kernel. As shown in Figure 5, the data (output) written by the kernel is the second argument. Since the OS knows all the buffers allocated by the process by intercepting the GPU memory allocation calls, line 1 (cudaMalloc), and it knows each kernel’s function signature [1], we can systematically compare the arguments with the allocated buffers to speculate the buffer written by each kernel.
Specifically, our speculation traces the kernel’s writes at buffer-level: When applications launch an opaque GPU computation kernel, for each argument, PhOS treats it as a tentative address pointing to a GPU buffer to write. To decide which buffer is written, we compare it with all the process’s allocated buffers. If the integer value of the argument falls within the range of a buffer, we mark the entire buffer as written. To reduce false positives, we filter out irrelevant arguments with the function signature. Specifically, we use clang [13] to extract the kernel’s argument types, focusing solely on mutable pointer arguments. One type that cannot be precisely filtered out is the C struct type, which is opaque to PhOS as we don’t know the struct definition. For such a type, we conservatively treat all 8-byte chunks in the struct as potential written GPU buffers.
Figure 6: PhOS validation workflow.
@@
Figure 6 presents the validation workflow. When PhOS encounters an opaque kernel that has not been instrumented (including JIT-compiled ones [58]), it generates a new twin kernel whose kernel binary is the instrumented version of the original kernel binary (❶). The instrumented validator performs the following: For each write instruction, it inserts address range checks before it to verify target address belongs to speculated buffers. If the validation fails, the validator reports the incident to PhOS by writing the address to a pre-allocated PhOS managed CPU buffer. The instrumentation is at the PTX ISA [51]-level for portability. The instrumentation overhead is negligible because for each kernel binary, it will be instrumented only once.
@@
Discussion: the granularity of tracing. Ideally, the trace should be as fine-grained as possible to avoid protocol overhead for ensuring correctness, e.g., avoid excessive copy-on-write. For non-opaque kernels, we can precisely track the bytes written. However, our speculation-based tracing can only detect writes at the buffer-level, which may result in over-tracing. For example, when a kernel only writes a small part of the buffer, we may treat the entire buffer as written. Luckily, such over-tracing is rare especially for recent GPU applications like AI jobs, because (1) AI frameworks (e.g., PyTorch) use a fine-grained buffer allocation, e.g., one for each tensor [59], and (2) the opaque kernels typically write the entire buffer (tensor). See §8.4 for the analysis.
@@
Suppose a checkpoint starts at time t1. Our CoW protocol guarantees the final checkpoint matches a stop-the-world checkpoint at t1, while allowing concurrent application execution. At a high level, the inconsistency can only happen during execution if the GPU writes to an uncheckpointed buffer. Before we allow such writes to execute, we first copy the data to an on-device buffer, then redirect application writes to a new buffer. Thus, the concurrent checkpoint will not see the new writes.
@@
Figure 7: An illustration of the soft copy-on-write protocol for concurrent checkpoint correctness.
@@
The protocol. Figure 7 presents the detailed protocol that consists of two phases: quiesce (➀) and concurrent copy (➁). The quiesce phase stops CPU and GPU execution, the same as existing stop-the-world checkpointing [63, 56]. Note that if we want to checkpoint a multi-process application, the quiesce phase stops all the processes’ CPUs and GPUs. Quiescing is necessary because it regulates the current process states to be the same as a possible stop-the-world checkpoint. To quiesce, we first stop CPU to prevent sending new GPU APIs. Then we wait for pending GPU kernels and communications to complete (e.g., via cudaDeviceSynchronize). Note that the quiescing overhead is negligible since GPU operations occur at microsecond scales. Thus, despite lacking concurrency, its overhead is significantly lower than data copying (recall Figure 2).
@@
At the copy phase, we copy all the GPU and CPU data to the checkpoint. For CPU data, we follow existing works [73, 66] to isolate concurrent writes with OS’s copy-on-write. For GPU, we isolate writes at the GPU API-level: Before executing a GPU kernel (or API), we check whether it writes to a buffer that has not been checkpointed (or is being concurrently checkpointed) using the traced writes through argument-based speculation. If so (e.g., B2), we will first copy the victim buffer to a new buffer allocated on the GPU (B2’), and then redirect the kernel to write to the new buffer (B2’). If multiple concurrent kernels try to write B2, the first kernel will copy B2 to B2’, and all the others will wait for the copy to finish before executing on B2’.
@@
Kernel buffer redirection, on-device buffer management and handling mis-speculation. Three things need to be noted. First, redirecting kernel writes to the new buffer is not that trivial: For types 1–3 APIs, we can directly change the arguments for so; but for the opaque ones, we cannot because the arguments can be mis-speculated. For example, if the changed argument is not a buffer, the execution will corrupt both application states as well as the checkpoint. Hence, we leave the kernel argument unchanged. To ensure correctness, we copy the buffer to a new on-device buffer, and let the checkpoint copy from the new buffer. A problem is that if an opaque kernel tries to write a buffer that is being concurrently checkpointed, it must be stalled to wait for the checkpoint copy to finish. Since the checkpoint copy is slower than the on-device copy, it may incur extra stalls. Fortunately, we rarely meet such cases empirically due to the low probability of the incidents.
@@
Second, we only reserve a small GPU memory (2 GB) for the copy-on-write, so during CoW we may lack sufficient memory to do the copy. Under such circumstances, we block the kernel until free memory is available (K2 in Figure 7). PhOS does not require a large on-device buffer because once a copy-on-write is done, its original buffer can be released. PhOS achieves so by holistically managing all GPU memory management by extending the related GPU APIs.
@@
Figure 8: An illustration of the soft dirty buffer protocol for concurrent checkpoint correctness.
@@
Once all the data has been copied, we start another quiesce phase to stop the CPU and GPU execution (➂). Note that dirty buffers written during the quiesce phase will also be recorded. After the quiesce is done, we recopy all the GPU dirty buffers CPU dirty pages to the checkpoint (➃). The application is stopped to ensure correctness, but we can also iteratively do the concurrent recopy similar to CPU-based protocols [14].
@@
While concurrent checkpoint can effectively minimize the process stalls, efficiently realizing it requires addressing several technical issues.
Figure 9: An illustration of how to reduce dirty buffers generated by prioritized CPU and GPU concurrent checkpoint. Note that the bars only show the execution on CPU and GPU, not the time when the dirty buffers can be generated.
Reducing dirty buffers via coordinated GPU and CPU checkpoint.   In the soft recopy protocol, the timing of the copy is critical to reducing the number of dirty buffers: if the copy is performed after the kernel write, the buffer will not become dirty so it will not be recopied. This requires copying buffers that are unlikely to be written first.
In GPU processes, we found the data written by the CPU and GPU differ: GPU writes more data than CPU. Thus, it is beneficial to first copy the CPU data instead of copying the GPU data, as shown in Figure 9 (b). However, naively checkpointing the CPU data and GPU data without control cannot achieve the aforementioned goal, because CPU and GPU checkpoints share the checkpoint bandwidth and thus interfere with each other (see Figure 9 (a)). Thus, we implement a coordinated checkpoint in our checkpoint engine: we first copy the CPU data, then the GPU data.
Prioritized application PCIe transfer.   When we completed our first concurrent checkpoint implementation (for both protocols), we found the process still stalls during the checkpoint. By profiling the utilization of GPU hardware, we found it is because the checkpoint data copy has saturated the GPU PCIe transfer engines. Specifically, GPUs have a limited number of PCIe transfer engines [7] shared between PhOS and applications. Without control, the bulk checkpoint load could easily saturate the PCIe engine, causing application execution starvation if it needs to load data from PCIe.
To this end, we adopt two techniques to avoid application starvation—the goal is to prioritize application PCIe transfer over the checkpoint, observing that applications infrequently use the PCIe but such usages are on their critical path. First, we intercept all process’s PCIe-related API calls (e.g., cudaMemcpy) to detect application PCIe transfer. Second, we implemented a preemption mechanism by using chunk-based checkpoint data copy. Specifically, for each GPU buffer, we copy it to the checkpoint in small 4 MB chunks. After copying one chunk, we check whether there is ongoing or pending application transfer. If so, we pause the checkpoint copy and let the pending application transfers execute first.
@@
Figure 10: An illustration of the concurrent restore in PhOS.
Concurrent and correct restore. When restoring data from the checkpoint to the GPUs, PhOS allows processes execution concurrently when PhOS is copying data from the checkpoint to the GPU buffers. Figure 10 illustrates the execution flow. To ensure correctness, before executing a kernel (or API), we check whether the data it requires has been copied or not. If not, we pause the execution, copy the buffer containing the required data on demand, and then resume. In our example, K0 will wait for B1 to be copied. If the data has been copied, e.g., K1 and K2, kernels can execute while the restore is concurrently executing (copying B3 and B4). This effectively overlaps computation with data transfer, reducing the observable impact of data copy overhead analyzed in §2.3.
Recall from §4 that the checkpoint of PhOS is correct. Given this guarantee, the correctness of concurrent restore depends on accurately tracing execution’s read set, similar to prior CPU systems [72, 20, 69]. We describe it next.
@@
Handling mis-speculation. Like writes, we also instrument a validator to ensure correctness under speculation failures. If the speculation is wrong, the validator will notify PhOS to handle. On the other hand, the mis-speculation handling is more complex, because the kernel may have accessed inconsistent GPU states caused by a partially restored buffer. To ensure correctness, we must roll back the GPU states to a correct state. Our current solution rolls back to the initial state from the checkpoint and then performs a stop-of-the-world restore for liveness. We choose this simple solution because we have not met speculation failures in common GPU applications. While a more efficient solution could involve isolating partial updates using our soft copy-on-write protocol, we defer this optimization for future work.
@@
To address the issue, our key observation is that GPU execution is decoupled from the context it uses: a process can use any context capable of execution. Therefore, we pre-create a pool of common GPU contexts at PhOS. Specifically, we maintain the pool in the PhOS daemon—a long-running OS service that pre-creates CUDA and cuBLAS contexts with cuCtxCreate and cublasCreate at boot time. For multi-GPU applications, we also pre-create an NCCL group communicator that covers all GPUs connected via NVLink, which can accelerate establishing the NCCL group communicator with sub-topology using ncclCommSplit [47] efficiently. We don’t create communicators across machines because the ideal communicator may depend on the network topology. The context pool operates in a separate process space, so applications must use inter-process communication mechanisms to access it.
With the context pool in PhOS daemon, when a GPU process requests to create a context, we intercept the context creation API call and assigns a pre-existing context from the pool. We also track the mapping between the context and the GPU process, ensuring all subsequent GPU API calls from that process utilize the mapped context.
@@
Providing fault tolerance to GPU processes (Checkpoint mostly).   Distributed computing applications such as training are susceptible to GPU failures [26, 22, 63]. PhOS provides transparent fault tolerance through periodic GPU process checkpointing using our soft copy-on-write (CoW) protocol described in §4.2. The checkpoints are stored in fast fault-tolerant storage, such as replicated remote memory [71]. If a failure occurs, PhOS stops all GPU processes, and then restores them from the most recent checkpoint. The checkpoint frequency is determined by holistically considering both the checkpoint overhead as well as the loss of computation due to restoring from a stale checkpoint, similar to prior work [28]. The detailed method is left in the appendix (§A.1). For fault tolerance, using a slightly stale checkpoint (i.e., one that does not capture the absolute latest execution state) is acceptable, because the latest state can be recovered through recomputation after restoring.
Two things to note. First, fault tolerance is a checkpoint-mostly case, as the checkpoint frequency is high to avoid losing computation results [71]. In comparison, the restore only happens upon failures, which occur much less frequently [46, 78]. Second, we need to ensure the checkpoint from all the involved processes is consistent [63]. Thus, we extended the quiescing phase described in §4.2 across all involved processes. After the quiesce, we can checkpoint each process with CoW separately.
Table 2: Applications evaluated. indicates a multi-GPU setup. PPO is training-only, and our setup cannot run Llama-3-70B training.
ResNET-${}^{\text{152M}}$) PPO-${}^{\text{336M}}$) Stable Diffusion-${}^{\text{1B}}$) Llama2-${}^{\text{13B}}$) Llama3.3-${}^{\text{70B}}$)
@@
Library Torchvision [60] OpenAI Gym [57] HuggingFace [24] Meta Llama 2 [43] Meta Llama 3.3 [44]
Figure 11: The application stall time caused by different OS-level checkpoint and restore systems.
Live migration of a GPU process (Checkpoint + Restore).   PhOS implemented pre-copy-style live migration [14]: It first checkpoints the process via our recopy protocol, then restores it on the target node. The recopy protocol is necessary because the destination should resume exactly at the last execution state at the source node. To avoid redundant data copy from first copying the data from source GPU to the checkpoint, then from the checkpoint to the target GPU, we use GPU-direct RDMA [2] to directly copy the data from the source GPUs’ buffers to the target GPUs’ buffers.
Fast GPU serverless function startups (Restore mostly).   In serverless computing, processes are created on demand for each request [33, 16]. Since each process has an entry function to react to a request, we take a checkpoint before the entry. Afterward, once a request comes, we restore the process from the checkpoint using our concurrent restore protocol described in §6. This eliminates the overhead to run till the entry function (commonly termed coldstart costs) [72, 5].
@@
We have implemented PhOS in 54,839 LoC in C++ and Rust, excluding the CUDA GPU driver extension framework, communication libraries and integration with CRIU [18].
Testbed. We conducted our experiments on two servers, each with eight NVIDIA A800 (80 GB HBM and 400 Gbps NVLink interconnects) GPUs, two Intel Xeon Gold 6348 CPUs (total 56 cores), and 1 TB of DRAM. Two servers are connected via an RDMA network with 100 Gbps bandwidth between each GPU and install CUDA 12.3 [48].
Baselines. We compare PhOS with NVIDIA’s official OS-level checkpoint and restore tool—cuda-checkpoint [56]. Nevertheless, we found it is extremely slow, e.g., it cannot achieve a PCIe-fully utilized data copy speed, see Figure 11. Yet, we don’t have its source code to analyze. To compare with the best performance of the stop-the-world C/R, we implemented Singularity [63]—the state-of-the-art stop-the-world GPU C/R system, in our codebase as the baseline. We have carefully tuned the implementation, e.g., we leverage pinned memory to achieve maximum data copy performance, to ensure a fair comparison with PhOS. As shown in Figure 11, our implementation of Singularity has orders of magnitude smaller application stall time than cuda-checkpoint. Thus, we will omit a detailed comparison with cuda-checkpoint in later analysis. Finally, a recent work CRIUgpu [64] builds on cuda-checkpoint to support a unified stop-the-world checkpoint and restore. Its objective is for functionalities but not performance. Hence, we omit a comparison because it has the same performance as cuda-checkpoint.
Figure 12: The average wasted GPU time (normalized) using different checkpoint methods for fault tolerance on training workloads. The wasted time is normalized to the total time used for training ($T$), and the displayed time is normalized to the number of systems with the maximum wasted time for ease of interpretation. cuda-checkpoint does not support checkpointing distributed jobs.
Figure 13: The downtime of migrating processes between machines with different systems. Multi-GPU workloads are marked with .
@@
We choose the end-to-end C/R workloads described in §7 to evaluate the effectiveness of PhOS.
@@
Fault tolerance: evaluated metrics and results.   To evaluate the fault tolerance performance of PhOS, we choose training applications and compare the performance of PhOS with baselines. Following existing works [71, 28], we compare the following metrics: checkpoint overhead and wasted GPU time during training. The checkpoint overhead measures the stall time caused by checkpointing, while the wasted GPU time is measured by the end-to-end training GPU time wasted when compared to an unrealistic, non-fault execution. The checkpoint is stored in host memory to avoid slow storage [71].

### v3

@@ Introduction
  3. (3)
    Issue computation kernels with well-defined semantics, e.g., cublasSgemm(...,A,B,...,C,...). They perform a computation and update the result buffers (C), whose read and write semantics are know in the specifications (cuBLAS, ).
  4. (4)
    Issue opaque computation kernels, e.g., cudaLaunch -Kernel (func, ... , args, ...) . They launch kernels written by the developers or just-in-time compiled during runtime, so the reads and writes are unknown to the OS.
@@
Figure 5. An illustration of the GPU kernel (fasttransformercode, ) written by the user (a) and how PhOS speculates the accessed buffers (b).
Opaque kernels are trickier because the OS only knows the kernel binary and the arguments invoking it. Our observation is that each GPU kernel represents a clear and straightforward computational purpose whose accessed data addresses are directly encoded in the OS-known arguments calling the kernel. As shown in Figure 5, the data (output) written by the kernel is the second argument. Since the OS knows all the buffers allocated by the process by intercepting the GPU memory allocation calls, line 1 (cudaMalloc), and it knows each kernel’s function signature (cuda-signature, ), we can systematically compare the arguments with the allocated buffers to speculate the buffer written by each kernel. The intercepting overhead is negligible: it only executes a few data structures manipulations atop of existing GPU memory allocation APIs.
Specifically, our speculation traces the kernel’s writes at buffer-level: When applications launch an opaque GPU computation kernel, for each argument, PhOS treats it as a tentative address pointing to a GPU buffer to write. To decide which buffer is written, we compare it with all the process’s allocated buffers. If the integer value of the argument falls within the range of a buffer, we mark the entire buffer as written. To reduce false positives, we filter out irrelevant arguments with the function signature. Specifically, we use clang (clang, ) to extract the kernel’s argument types, focusing solely on mutable pointer arguments. One type that cannot be precisely filtered out is the C struct type, which is opaque to PhOS as we don’t know the struct definition. For such a type, we conservatively treat all 8-byte chunks in the struct as potential written GPU buffers.
Figure 6. PhOS validation workflow.
@@
Figure 6 presents the validation workflow. When PhOS encounters an opaque kernel that has not been instrumented (including JIT-compiled ones (torch-jit, )), it generates a new twin kernel whose kernel binary is the instrumented version of the original kernel binary (❶). The instrumented validator performs the following: For each write instruction, it inserts address range checks before it to verify target address belongs to speculated buffers. If the validation fails, the validator reports the incident to PhOS by writing the address to a pre-allocated PhOS managed CPU buffer. The instrumentation is at the PTX ISA (ptx, )-level for portability. The instrumentation overhead is negligible because for each kernel binary, it will be instrumented only once.
@@
Discussion: the granularity of tracing.    Ideally, the trace should be as fine-grained as possible to avoid protocol overhead for ensuring correctness, e.g., avoid excessive copy-on-write. For non-opaque kernels, we can precisely track the bytes written. However, our speculation-based tracing can only detect writes at the buffer-level, which may result in over-tracing. For example, when a kernel only writes a small part of the buffer, we may treat the entire buffer as written. Luckily, such over-tracing is rare especially for recent GPU applications like AI jobs, because (1) AI frameworks (e.g., PyTorch) use a fine-grained buffer allocation, e.g., one for each tensor (pytorch-mem, ), and (2) the opaque kernels typically write the entire buffer (tensor). See §8.5 for the analysis.
@@
Suppose a checkpoint starts at time t1. Our CoW protocol guarantees the final checkpoint matches a stop-the-world checkpoint at t1, while allowing concurrent application execution. At a high level, the inconsistency can only happen during execution if the GPU writes to an uncheckpointed buffer. To fix the issue, before we allow such writes to execute, we first copy the targeted data to another place, e.g., a free on-device buffer; then redirect application writes to a new buffer. Thus, the concurrent checkpoint will not see the new writes.
@@
Figure 7. An illustration of the soft copy-on-write protocol for concurrent checkpoint correctness.
@@
The protocol.   Figure 7 presents the detailed protocol that consists of two phases: quiesce (➀) and concurrent copy (➁). The quiesce phase stops CPU and GPU execution, the same as existing stop-the-world checkpointing (DBLP:journals/corr/abs-2202-07848, ; nvidia-new-ckpt-driver-repo, ). Note that if we want to checkpoint a multi-process application, the quiesce phase stops all the processes’ CPUs and GPUs. Quiescing is necessary because it regulates the current process states to be the same as a possible stop-the-world checkpoint. To quiesce, we first stop CPU to prevent sending new GPU APIs. Then we wait for pending GPU kernels and communications to complete (e.g., via cudaDeviceSynchronize). Note that the quiescing overhead is negligible since GPU operations occur at microsecond scales. Thus, despite lacking concurrency, its overhead is significantly lower than data copying (recall Figure 2).
@@
At the copy phase, we copy all the GPU and CPU data to the checkpoint. For CPU data, we follow existing works (DBLP:conf/sosp/Wu0M023, ; Aurora, ) to isolate concurrent writes with OS’s copy-on-write. For GPU, we isolate writes at the GPU API-level: Before executing a GPU kernel (or API), we check whether it writes to a buffer that has not been checkpointed (or is being concurrently checkpointed) using the traced writes through argument-based speculation. If so (e.g., B2), we will first copy the victim buffer to a new buffer allocated on the GPU (B2’), and then redirect the kernel to write to the new buffer (B2’). If multiple concurrent kernels try to write B2, the first kernel will copy B2 to B2’, and all the others will wait for the copy to finish before executing on B2’. If the GPU has no free memory to allocate the new buffer, we will directly copy the data to the checkpoint at the host.
@@
Kernel buffer redirection, on-device memory used, and handling mis-speculation.    Three things need to be noted. First, redirecting kernel writes to the new buffer is not that trivial: For types 1–3 APIs, we can directly change the arguments for so; but for the opaque ones, we cannot because the arguments can be mis-speculated. For example, if the changed argument is not a buffer, the execution will corrupt both application states as well as the checkpoint. Hence, unlike traditional CPU copy-on-write, we leave the kernel argument unchanged, and redirect the checkpoint to access the copied buffer instead.
@@
Second, we only reserve a small GPU memory (up to 2 GB) for the copy-on-write, and the concurrent checkpoint can still work even if no extra memory is used. If we lack sufficient GPU memory for the copy-on-write, we block the kernel until free memory is available (K2 in Figure 7) or the written buffer has been checkpointed. This causes a little checkpoint stall but we found it acceptable: PhOS does not require a large on-device buffer for high performance because once a copy-on-write is done, its original buffer can be immediately released. §8.3 analyzes the impact of GPU memory reserved for CoW on the checkpointing performance.
@@
Discussion: speculation vs. validation.    Although we can, in principle, avoid speculation by instrumenting the kernels to obtain their accessed buffers and do the copy-on-write, as we did in validation, we still choose speculation for two reasons. First, speculation allows us to acquire buffer information in advance, which is necessary for CoW especially in case of insufficient on-device buffer. Besides, our concurrent restore protocol (see §6) also relies on knowing the accessed buffers in advance. Second, speculation then validation has fewer overheads than obtaining the buffers without speculation, because in the common case, the validation will only pass a flag indicating whether the validation passes.
@@
Figure 8. An illustration of the soft dirty buffer protocol for concurrent checkpoint correctness.
@@
Once all the data has been copied, we start another quiesce phase to stop the CPU and GPU execution (➂). Note that dirty buffers written during the quiesce phase will also be recorded. After the quiesce is done, we recopy all the GPU dirty buffers CPU dirty pages to the checkpoint (➃). The application is stopped to ensure correctness, but we can also iteratively do the concurrent recopy similar to CPU-based protocols (DBLP:conf/nsdi/ClarkFHHJLPW05, ).
@@
Figure 9. An illustration of the reduced time window size that could cause dirty buffers thanks to prioritized CPU and GPU concurrent checkpoint. The bars show the time window size.
Reducing dirty buffers via coordinated GPU and CPU checkpoint.    In the soft recopy protocol, the time window of the copy is critical to the number of dirty buffers generated: the longer the copy, the more likely the buffers are becoming dirty. If we copy the CPU and GPU data without coordination, the overall time window is the size of checkpoint image divided by the copy bandwidth.
In GPU processes, we found GPU writes data more frequently than CPU writes, so it is beneficial to first copy the CPU data instead of copying the GPU data, reducing the GPU time window to the size of the GPU data divided by the copy bandwidth, as shown in Figure 9 (b). Thus, we implement a coordinated checkpoint in our checkpoint engine: we first copy the CPU data, then the GPU data.
Prioritized application PCIe transfer.    When we completed our first concurrent checkpoint implementation (for both protocols), we found the process still meets significant stalls during the checkpoint. By profiling the GPU execution, we found the stall is not caused by our protocols, but by the saturated DMA engine caused by GPU checkpoint. Specifically, GPUs have a limited number of DMA engines (bakitademystifying, ) shared between PhOS and applications. With concurrent checkpoint, the bulk copy nature of the checkpoint can easily saturate the DMA engines, causing application kernels to wait.
To this end, we adopt two techniques to avoid application starvation—the goal is to prioritize application DMA transfer over the checkpoint, observing that applications infrequently use the DMA but such usages are on their critical path. First, we intercept all process’s DMA-related API calls (e.g., cudaMemcpy) to detect application PCIe transfer. The interception overhead is negligible (<1%), because the APIs are typically called asynchronously so our interception is not on the application’s critical path (DBLP:journals/corr/abs-2401-13354, ; DBLP:journals/corr/abs-2306-03622, ). Based on the detection, we implemented a preemptible checkpoint copy mechanism for the application data copy: for each GPU buffer, we copy it to the checkpoint in small 4 MB chunks. After copying one chunk, we check whether there is ongoing or pending application transfer. If so, we pause the checkpoint copy and let the pending application transfers execute first.
@@
Figure 10. An illustration of the concurrent restore in PhOS.
Concurrent and correct restore.    When restoring data from the checkpoint to the GPUs, PhOS allows processes to execute concurrently when PhOS is copying data from the checkpoint to the GPU buffers. Figure 10 illustrates the execution flow. To ensure correctness, before executing a kernel (or API), we check whether the data it requires has been copied. If not, we pause the execution, copy the buffer containing the required data, and then resume. In our example, K0 will wait for B1 to be copied. If the data has been copied, e.g., K1 and K2, kernels can execute while the restore is concurrently executing (copying B3 and B4). This effectively overlaps computation with data transfer, reducing the observable impact of data copy overhead analyzed in §2.3.
Recall from §4 that the checkpoint of PhOS is correct. Given this guarantee, the correctness of concurrent restore depends on accurately tracing execution’s read set, similar to prior CPU systems (DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/asplos/DuYXZYQWC20, ; DBLP:conf/eurosys/WangHW19, ).
@@
Handling mis-speculation.    Like writes, we also instrument a validator to ensure correctness under speculation failures. If the speculation is wrong, the validator will notify PhOS to handle. However, the mis-speculation handling of reads is more complex than writes, because the kernel may have accessed inconsistent GPU states caused by a partially restored buffer. To ensure correctness, we must roll back the GPU states to a correct one. Our current solution rolls back to the initial state from the checkpoint and then performs a stop-of-the-world restore for liveness. We choose this simple solution because we have not met speculation failures in common GPU applications. While a more efficient solution could involve isolating partial updates using our soft copy-on-write protocol, we defer the detailed design to future work.
@@
To address the issue, our key observation is that GPU execution can be decoupled from the context it uses: a process can use any context capable of executing the APIs. Therefore, we pre-create a pool of common GPU contexts at PhOS. Specifically, we maintain the pool in the PhOS daemon—a long-running OS service that pre-creates CUDA and cuBLAS contexts with cuCtxCreate and cublasCreate at boot time. For multi-GPU applications, we also pre-create an NCCL group communicator that covers all GPUs connected via NVLink, which can accelerate establishing the NCCL group communicator with sub-topology using ncclCommSplit (ncclgroup, ) efficiently. We don’t create communicators across machines because the ideal communicator may depend on the network topology. The context pool operates in a separate process space, so applications must use inter-process communication mechanisms to access it.
With the context pool, when a GPU process requests to create a context, we intercept the context creation API call and assign a pre-existing context from the pool. We also track the mapping between the context and the GPU process, ensuring all subsequent GPU API calls from that process utilize the mapped context.
@@
Fault tolerance for GPU processes (checkpoint-mostly).    Distributed computing applications such as training models are susceptible to GPU failures (DBLP:conf/icse/GaoSLZWLY23, ; DBLP:conf/nsdi/EisenmanMIMKNSA22, ; DBLP:journals/corr/abs-2202-07848, ). PhOS provides transparent fault tolerance through periodic GPU process checkpointing using our soft copy-on-write (CoW) protocol described in §4.2. The checkpoints are stored in fast fault-tolerant storage, such as with replications stored in remote memory (DBLP:conf/sosp/WangJZZFNW23, ). If a failure occurs, PhOS first stops all GPU processes, and then restores them from the most recent checkpoint. The checkpoint frequency is determined by holistically considering both the checkpoint overhead as well as the loss of computation due to restoring from a stale checkpoint, similar to prior work (DBLP:conf/eurosys/GuptaKKVGKRS24, ). The detailed method is left in the appendix (§A.1). For fault tolerance, using a slightly stale checkpoint (i.e., one that does not capture the absolute latest execution state) is acceptable, because the latest state can be recovered through recomputation.
Two things to note. First, fault tolerance is a checkpoint-mostly case, as the checkpoint frequency needs to be high to avoid losing computation results (DBLP:conf/sosp/WangJZZFNW23, ). In comparison, the restore only happens upon failures, which occur much less frequently, e.g., one per-hour (DBLP:conf/fast/MohanPC21, ; DBLP:journals/corr/abs-2205-01068, ). Second, we need to ensure the checkpoint from all the involved processes is consistent (DBLP:journals/corr/abs-2202-07848, ). Thus, we extended the quiescing phase described in §4.2 across all involved processes. After the quiesce, we can checkpoint each process with CoW separately. We currently follow Singularity (DBLP:journals/corr/abs-2202-07848, ) and use a user-provided hint for a correct global quiesce.
Table 2. Applications evaluated. indicates a multi-GPU setup. PPO is training-only, and our setup cannot run Llama3.3-70B training.
[…]
@@
Library Torchvision (torchvision, ) OpenAI Gym (gym, ) HuggingFace (huggingface, ) Meta Llama 2 (llama, ) Meta Llama 3.3 (llama3, )
Figure 11. The application stall time caused by different OS-level checkpoint and restore systems.
Live migration of a GPU process (checkpoint-restore).  PhOS implemented pre-copy-style live migration (DBLP:conf/nsdi/ClarkFHHJLPW05, ): It first checkpoints the process via our recopy protocol, then restores it on the target node. The recopy protocol is necessary because the destination should resume exactly at the last execution state at the source node. To avoid redundant data copy from first copying the data from source GPU to the checkpoint, then from the checkpoint to the target GPU, we use GPU-direct RDMA (gpu-direct-rdma, ) to directly copy the data from the source GPUs’ buffers to the target GPUs’ buffers.
Fast GPU serverless function startups (restore-mostly).    In serverless computing, processes are created on demand for each request (DBLP:journals/corr/abs-1902-03383, ; ali-serverless-gpu, ). Since each process has an entry function to react to a request, we take a checkpoint right before the entry, so once a request comes, we can restore the process from the checkpoint using our concurrent restore protocol described in §6 to bypass the cold start cost (DBLP:conf/osdi/WeiLW0Y0023, ; DBLP:conf/eurosys/AoPV22, ) before the entry.
@@
We have implemented PhOS in 54,839 LoC in C++ and Rust, excluding the CUDA GPU driver extension framework, communication libraries and integration with CRIU (criu, ).
Testbed.    We conducted our experiments on A800 servers, each with eight NVIDIA A800 (80 GB HBM and 400 Gbps NVLink interconnects) GPUs, two Intel Xeon Gold 6348 CPUs (total 56 cores), and 1 TB of DRAM. Servers are connected via an RDMA network with 100 Gbps bandwidth between each GPU and install CUDA 11.3 (cuda123, ).
Baselines.   We compare PhOS with NVIDIA’s official OS-level checkpoint and restore tool—cuda-checkpoint (nvidia-new-ckpt-driver-repo, ), which adopts a stop-the-world approach. Nevertheless, we found it is extremely slow, e.g., it cannot fully utilize the PCIe bandwidth to copy data, see Figure 11. Unfortunately, we don’t have its source code to analyze. To compare with the best performance of the stop-the-world C/R, we implemented Singularity (DBLP:journals/corr/abs-2202-07848, )—the state-of-the-art GPU C/R system, in our codebase as the baseline. We have carefully tuned the implementation, e.g., we leverage pinned memory to achieve maximum data copy performance. As shown in Figure 11, our implementation of Singularity has orders of magnitude smaller application stall time than cuda-checkpoint. Thus, we will omit a detailed comparison with cuda-checkpoint in later analysis. Finally, a recent work CRIUgpu (DBLP:journals/corr/abs-2502-16631, ) builds on cuda-checkpoint to support a unified stop-the-world checkpoint and restore. Its objective is for functionalities but not performance. Hence, we omit a comparison because it has the same performance as cuda-checkpoint.
Figure 12. The wasted GPU time using different checkpoint methods for fault tolerance on training workloads. The reported wasted time is normalized to the total time used for training ($T$), and the displayed time is normalized to the number of systems with the maximum wasted time for ease of interpretation. cuda-checkpoint does not support checkpointing distributed jobs.
Figure 13. The downtime of migrating processes between machines with different systems. Multi-GPU workloads are marked with .
@@
Fault tolerance: evaluated metrics and results.    To evaluate the effectiveness of PhOS in fault tolerance scenarios, we choose training applications. Following existing works (DBLP:conf/sosp/WangJZZFNW23, ; DBLP:conf/eurosys/GuptaKKVGKRS24, ), we compare different systems with the following metrics: checkpoint overhead and wasted GPU time during training. The checkpoint overhead measures the stall time caused by checkpointing, while the wasted GPU time measures the end-to-end training GPU time wasted when compared to an unrealistic, non-fault execution. For all systems, the checkpoint is stored in host memory to avoid slow storage (DBLP:conf/sosp/WangJZZFNW23, ).